# Gravitational waves from vacuum bubbles: ultraviolet dependence on wall thickness

### v1

@@ VI Numerical fitting
To first determine the values of $l_{\rm col}$, we should obtain the bubble collosion time. With the bubble profiles simulated in Sec. IV, we can extract the time evolution of the bubble radius $R(t)$ and wall thickness $l(t)$. Then, compare $R(t)$ to the half of the bubble separation, $d/2=R_{\rm col}$, the bubble collision time $t_{\rm col}$ is thus determined. Finally we obtain $l_{\rm col}\equiv l(t_{\rm col})$ and hence the ratio $l_{\rm col}/R_{\rm col}\equiv l(t_{\rm col})/R(t_{\rm col})$ at bubble collision. With all preparations completed, we numerically fit the relationship between the high-frequency power and $l_{\rm col}/R_{\rm col}$. The results are shown in Fig. 6. We observe that the absolute value of the power decreases as $l_{\rm col}/R_{\rm col}$ increases, which can be perfectly fitted with a single power law as
@@ VII Analytical approximation
In Sec. VI, we have derived the fitting results for the relationship between the high-frequency power and the wall thickness (normalized to the bubble radius) $l_{\rm col}/R_{\rm col}$ at collisions, as given in (31)-(33). Although an immediate understanding of this result is currently not feasible, for specific FOPTs originating from different new physics models, calculating $l_{\rm col}/R_{\rm col}$ directly in terms of other more model-independent characteristics is still of large theoretical interest.
We estimate the value of $l_{\rm col}/R_{\rm col}$ from the energy conservation. First, we assume that before and at the moment of the bubble collision, the interaction between bubbles is negligible. Thus, the total energy is simply the sum of the energies of individual bubbles, each of which is conserved. Here, we make the following ansatz for the scalar profile $\phi$ of a single bubble,
@@
where $v_{w}(t)\equiv\mathrm{d}R(t)/\mathrm{d}t$ is the bubble wall velocity. The reason for neglecting the second term in the square brackets is as follows: when integrating $(\partial_{t}\phi)^{2}$ to get $E_{\rm K}$, the dominant contribution comes from the region where $r\sim R(t)$, and the term $r-R(t)$ is nearly zero in this region. Then the kinetic energy at collision (22) can be approximated as
@@
$E_{V}(t_{\rm col})\sim-\frac{4\pi}{3}R_{\rm col}^{3}\Delta V+4\pi R_{\rm col}^{2}l_{\rm col}V_{b}.$ (41)
The first term in (41) corresponds to the contribution inside the bubble, while the second term represents the contribution at the bubble wall. Our simulation results demonstrate that $R(t)$ increases, while $l(t)$ decreases as time progresses. When $t_{\rm col}$ is relatively large, the second term in (41) becomes much smaller than the first term. Therefore, we have
$E_{V}(t_{\rm col})\sim-\frac{4\pi}{3}R_{\rm col}^{3}\Delta V.$ (42)
@@
We further simplify (44) by neglecting $E_{T}$. This is because that, for the quantum phase transition, $E_{T}$ is simply zero, while for the semi-classical phase transition, the absolute value of the potential energy significantly exceeds the total energy as shown in Fig. 3 as the time evolves. Thus, we can safely approximate $E_{T}\ll 4\pi R^{3}_{\rm col}\Delta V$ and finally, Eq. (44) reduces to
@@ VIII Conclusions and discussions
Investigating the stochastic GW signals offers a promising way to probe cosmological FOPTs and underlying new physics models. To gain a deeper understanding of the FOPT, it is crucial to analyze specific characteristics of the GW spectrum, such as high-frequency power law. In this work, we study the relationship between high-frequency power law and the ratio of bubble wall thickness to bubble radius at collision using numerical simulations and express this ratio as a function of model-independent parameters. Note that, to prepare the initial configurations for the numerical simulation, we derive for the first time the general expressions for the bubble number density and mean bubble separation in the vacuum first-order phase transition for runaway bubbles instead of a constant wall velocity. Additionally, to provide a more comprehensive analysis, we also consider the semi-classical phase transition and calculate for the first time the associated GWs.
Although we have built the relationship (31) between the high-frequency power law and the bubble-wall thickness at collisions, the physical explanation for such a relation, along with the observed negative correlation, remains unclear. In the future, we aim to derive this relationship analytically, which will enhance our understanding of the dynamics of cosmological FOPTs. Furthermore, larger numerical simulations for collisions of more bubbles are also necessary to check our preliminary results to greater generality, though we expect the qualitative picture may remain the same. Finally, more general simulations should also be investigated for the thermal phase transition with the exponent in the decay rate beyond the linear order and with the given backreaction force beyond the leading order. In this case, the difficulty lies in the clean extraction from the total GWs for the pure scalar contribution of wall collisions, which is nevertheless sub-dominated to the sound waves.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China Grants No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12105344, No. 11821505, No. 12235019, No. 11991052, No. 12047503, No. 12073088, and No. 11947302, the Science Research Grants from the China Manned Space Project with No. CMS-CSST-2021-B01 (supported by China Manned Space Program through its Space Application System).

### v3

@@ VI Numerical fitting
To first determine the values of $l_{\rm col}$, we should obtain the bubble collision time. With the bubble profiles simulated in Sec. IV, we can extract the time evolution of the bubble radius $R(t)$ and wall thickness $l(t)$. Then, compare $R(t)$ to the half of the bubble separation, $d/2=R_{\rm col}$, the bubble collision time $t_{\rm col}$ is thus determined. Finally we obtain $l_{\rm col}\equiv l(t_{\rm col})$ and hence the ratio $l_{\rm col}/R_{\rm col}\equiv l(t_{\rm col})/R(t_{\rm col})$ at bubble collision. With all preparations completed, we numerically fit the relationship between the high-frequency power and $l_{\rm col}/R_{\rm col}$. The results are shown in Fig. 6. We observe that the absolute value of the power decreases as $l_{\rm col}/R_{\rm col}$ increases, which can be perfectly fitted with a single power law as
@@
Finally, it is important to note that Ref. Gould et al. (2021) performs a similar numerical simulation for a two-bubble system, investigating the dependence of the high-frequency power $a_{n}$ on the scalar-field potential parameter and the bubble-wall Lorentz factor at collision time. The authors demonstrate that the difference in $a_{n}$ derived through various potential parameters becomes less pronounced at larger Lorentz factors. In contrast, in our study, once the scalar potential parameters are fixed, the mean bubble separation $d$ is also fixed, thereby determining the bubble Lorentz factor at collision time as a derived quantity rather than an independent parameter. Consequently, the explicit dependence of $a_{n}$ on the Lorentz factor is not apparent in our fitting result given by (31).
@@ VII Analytical approximation
In Sec. VI, we have derived the fitting results for the relationship between the high-frequency power and the wall thickness (normalized to the bubble radius) $l_{\rm col}/R_{\rm col}$ at collisions, as given in (31)-(33). Although an immediate understanding of this result is currently not feasible, for specific FOPTs originating from different new physics models, calculating $l_{\rm col}/R_{\rm col}$ directly in terms of other more model-independent characteristics is still of large theoretical interest. Such calculations could help distinguish between various new physics models that lead to a first order phase transition.
We estimate the value of $l_{\rm col}/R_{\rm col}$ from the energy conservation. The basic idea is that we first express the kinetic, gradient, and potential energies of a single bubble in terms of $l_{\rm col}$, $R_{\rm col}$, and other model-independent characteristics. Then by applying the energy conservation condition, we can obtain approximate expressions of $l_{\rm col}/R_{\rm col}$. Specifically, we first assume that before and at the moment of the bubble collision, the interaction between bubbles is negligible. Thus, the total energy is simply the sum of the energies of individual bubbles, each of which is conserved. Here, we make the following ansatz for the scalar profile $\phi$ of a single bubble,
@@
[…]
@@
$E_{V}(t_{\rm col})=\int_{0}^{\infty}\mathrm{d}r4\pi r^{2}V(\phi(r,t_{\rm col}))$ (41)
$\simeq\int_{0}^{R_{\rm col}}\mathrm{d}r4\pi r^{2}V(\phi_{-})+\int_{R+l_{\rm col}}^{\infty}\mathrm{d}r4\pi r^{2}V(\phi_{+})$
$+4\pi R_{\rm col}^{2}\int_{R_{\rm col}}^{R_{\rm col}+l_{\rm col}}\mathrm{d}rV(\phi(r,t_{\rm col}))$
$\simeq-\frac{4\pi}{3}R_{\rm col}^{3}\Delta V+4\pi R_{\rm col}^{2}\sigma^{V}_{\rm col}.$
Here, the second equality indicates that we have divided $E_{V}(t_{\rm col})$ into contributions from inside the bubble, outside the bubble, and the bubble surface region and have employed the approximation $l_{\rm col}\ll R_{\rm col}$. Furthermore, in deriving the third equality, we have used the conditions $V(\phi_{+})=0$ and the definition $\Delta V=V(\phi_{+})-V(\phi_{-})$. In the last line of (41), we have separated out the potential contribution to the bubble surface tension at the collision time, $\sigma^{V}_{\rm col}$, explicitly given by
$\sigma^{V}_{\rm col}=\int_{R_{\rm col}}^{R_{\rm col}+l_{\rm col}}\mathrm{d}rV(\phi(r,t_{\rm col})).$ (42)
It is important to note that this contribution is part of the total contribution of the bubble surface tension Linde (1983) that typically includes both potential and gradient energy contributions, i.e., $\sigma=\int\mathrm{d}r[(\mathrm{d}\phi/\mathrm{d}r)^{2}/2+V(\phi)]$. The numerical values of $\sigma^{V}_{\rm col}$ obtained from our simulations are listed in the last column of Table 2. Our numerical results indicate that the second term in the final line of Eq. (41) is significantly smaller than the first term. Thus, in subsequent analyses, we approximate $E_{V}(t_{\rm col})$ simply as $-4\pi R_{\rm col}^{3}\Delta V/3$.
@@
We further simplify (44) by neglecting $E_{T}$. This is because that, for the quantum phase transition, $E_{T}$ is simply zero, while for the semiclassical phase transition, the absolute value of the potential energy significantly exceeds the total energy as shown in Fig. 3 as the time evolves. Thus, we can safely approximate $E_{T}\ll 4\pi R^{3}_{\rm col}\Delta V$ and finally, Eq. (44) reduces to
@@
Finally, we emphasize that our derivation relies on four key assumptions: (1) strict energy conservation for isolated bubbles before collision; (2) adopting the simplified scalar field profile given in (34), valid mainly for relatively thin bubble walls; (3) bubble wall thickness being significantly smaller than bubble radius at collision; and (4) the total energy of each bubble being substantially smaller than the absolute potential energy magnitude. These assumptions collectively restrict our results to scenarios characterized by thin walls and large bubble separations. Therefore, our findings are limited and preliminary. In our future work, we will provide a more rigorous and generalized treatment of $l_{\rm col}/R_{\rm col}$ in terms of universally applicable model-independent characteristics.
@@ VIII Conclusions and discussions
Investigating the stochastic GW signals offers a promising way to probe cosmological FOPTs and underlying new physics models. To gain a deeper understanding of the FOPT, it is crucial to analyze specific characteristics of the GW spectrum, such as high-frequency power law. In this work, we study the relationship between high-frequency power law and the ratio of bubble wall thickness to bubble radius at collision using numerical simulations and express this ratio as a function of model-independent parameters. Note that, to prepare the initial configurations for the numerical simulation, we derive for the first time the general expressions for the bubble number density and mean bubble separation in the vacuum first-order phase transition for runaway bubbles instead of a constant wall velocity. Additionally, to provide a more comprehensive analysis, we also consider the semiclassical phase transition and calculate for the first time the associated GWs.
Although we have built the relationship (31) between the high-frequency power law and the bubble-wall thickness at collisions, the physical explanation for such a relation, along with the observed negative correlation, remains unclear. In the future, we aim to derive this relationship analytically, which will enhance our understanding of the dynamics of cosmological FOPTs. Furthermore, larger numerical simulations for collisions of more bubbles are also necessary to check our preliminary results to greater generality, though we expect the qualitative picture may remain the same. Finally, more general simulations should also be investigated for the thermal phase transition with the exponent in the decay rate beyond the linear order and with the given backreaction force beyond the leading order. In this case, the difficulty lies in the clean extraction from the total GWs for the pure scalar contribution of wall collisions, which is nevertheless subdominated to the sound waves.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China Grants No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12105344, No. 11821505, No. 12235019, No. 11991052, No. 12047503, No. 12073088, and No. 11947302, the Science Research Grants from the China Manned Space Project with No. CMS-CSST-2021-B01 and No. CMS-CSST-2025-A01. Our code and data have been made public on the website: https://github.com/Jun-Chen-Wang/Two-Bubble-Simulation/tree/main.